OpenStudy (anonymous):

@jim_thompson5910 I need you help wise one!! Jordan finances $459,600 with a 25/9 balloon mortgage at 6.9%. How much will he pay for principle and interest over the life of the loan? A) $491,312.40 B) $965,730.00 C) $721,306.34 D) $723,454.79

jimthompson5910 (jim_thompson5910):

sorry just noticed this now

jimthompson5910 (jim_thompson5910):

I would use that same calculator use it to determine his monthly payment and what the balloon balance will be

jimthompson5910 (jim_thompson5910):

let me know what you get

OpenStudy (anonymous):

what would I put in for the years??

jimthompson5910 (jim_thompson5910):

25 because of the "25/9"

jimthompson5910 (jim_thompson5910):

"25/9" means you have a 25 yr mortgage but the balloon payment hits right after year 9

OpenStudy (anonymous):

ok that helps a lot!

jimthompson5910 (jim_thompson5910):

what monthly payment do you get

OpenStudy (anonymous):

I got 3,219.10 but it's the same every month and I thought it was supposed to change

jimthompson5910 (jim_thompson5910):

you're thinking of an ARM

OpenStudy (anonymous):

Im so confused...

jimthompson5910 (jim_thompson5910):

ARM = adjustable rate mortgage

jimthompson5910 (jim_thompson5910):

what is the remaining balance in June 2024 ?

OpenStudy (anonymous):

$373,643.54

jimthompson5910 (jim_thompson5910):

you pay 9 years worth of monthly payments of $3,219.10 (3,219.10 per month)*(12 months per year)*(9 years) = 347,662.80

jimthompson5910 (jim_thompson5910):

Then you add on the remaining balance because that's the balloon payment

OpenStudy (anonymous):

and the remaining balance I would be on June of 2024 right?

jimthompson5910 (jim_thompson5910):

yeah the remaining balance there

OpenStudy (anonymous):

ok, the answer I got is $721,295.54 but that's not an option. it's close to one but not exact. now what?

jimthompson5910 (jim_thompson5910):

you pay 9 years worth of monthly payments of $3,219.10 (3,219.10 per month)*(12 months per year)*(9 years) = 347,662.80 then add on the remaining balance for june 2024 (principle+interest) + (balloon payment) = (347,662.80)+(373,643.54) = 721,306.34

OpenStudy (anonymous):

You are a life saver my friend!!! None of my teachers could have explained it to me the way you have! Thank You So Freaking Much!!! :D

jimthompson5910 (jim_thompson5910):

no problem